OpenStudy (anonymous):

Meats must be cooked at high enough temperatures to kill harmful bacteria. According to the USDA and the FDA, the internal temperature for cooked roasts and steaks should be at least 145°F, and whole poultry should be 180°F. Convert these temperatures to degrees Celsius. Round up to the next degree.

OpenStudy (anonymous):

cheat http://www.csgnetwork.com/tempconvjava.html

OpenStudy (anonymous):

formula is \[C=\frac{5}{9}(F-32)\]

OpenStudy (anonymous):

isnt the formula the other one

OpenStudy (anonymous):

C\[C=\frac{5}{9}(145-32)\] \[C=\frac{5}{9}(113)\] \[C=62.777\]

OpenStudy (anonymous):

the formula i wrote converts from F to C which is what you want to do. there is obviously a different one to convert from C to F

OpenStudy (anonymous):

ok cool i get confused by them

OpenStudy (anonymous):

its incorrect

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

let me try again.

OpenStudy (anonymous):

oh 63

OpenStudy (anonymous):

my fault sorry

OpenStudy (anonymous):

i need to do 180 also i just dnt now how to get the last part

OpenStudy (anonymous):

\[\frac{5}{9}(180-32)=\frac{5}{9}(148)=82.2\] rounded

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

actually i am certain it is right. what answer is it supposed to be? if you round up it is 83

OpenStudy (anonymous):

i had to round up thnk you

OpenStudy (anonymous):

do you now how to do this one Make the conversion. (Round your answer to one decimal place.) 5.8 oz = 1 g